A metal disc of diameter 20cm makes 1200rpm about its axis kept parallel to a uniform magnetic field of induction 10^-2T. The emf induced between the the centre and the rim of the disc is?

Grade:12th pass

2 Answers

PRASHANT KUMAR
5 years ago
Zero. Because flux is B.ds and here angle is 90 degree so its zero and emf is rate of change of flux so zero.
ketaka singh
5 years ago
Frequency of rotation = 20 cycles per sec implies \omega = 2\pi f = 40 \pi rad/sec
Consider a point A on the circumference if the disc.
Now consider a small element of length dx at a distance x from the centre if the disc on the line joining A and the centre.
 
Emf induced in this length dx , de = B(dx)v = B\omegaxdx 
Net emf induced between the centre and point A is 
 
E = \int de=\int_{x=0}^{x=r}B\omega xdx= B\omega r^{2}\frac{1}{2}
 
Calculating we get E = 6.28 × 10^-3